The Paulson Company's year-end balance sheet is shown below. Its cost of common equity is 16%, its before-tax cost of debt is 8%, and its marginal tax rate is 25%. Assume that the firm's long-term debt sells at par value. The firm’s total debt, which is the sum of the company’s short-term debt and long-term debt, equals $1,108. The firm has 576 shares of common stock outstanding that sell for $4.00 per share.

Assets Liabilities And Equity
Cash $ 120 Accounts payable and accruals $ 10
Accounts receivable 240 Short-term debt 58
Inventories 360 Long-term debt 1,050
Plant and equipment, net 2,160 Common equity 1,762
Total assets $2,880 Total liabilities and equity $2,880

Calculate Paulson's WACC using market-value weights. Do not round intermediate calculations. Round your answer to two decimal places
Answer #1

MV of equity = price*shares = 4*576 = 2304

Total Capital value = Value of Debt + Value of Equity
=1108+2304
=3412
Weight of Debt = Value of Debt/Total Capital Value
= 1108/3412
=0.3247
Weight of equity = 1-D/A
Weight of equity = 1-0.3247
W(E)=0.6753
After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 8*(1-0.25)
= 6
WACC=after tax cost of debt*W(D)+cost of equity*W(E)
WACC=6*0.3247+16*0.6753
WACC =12.75%